OpenStudy (anonymous):

Let f : Z × Z → Z × Z be defined as f(m, n) = (3m + 7n, 2m + 5n). Is f a bijection, i.e.,one- to-one and onto? If yes then give a formal proof, based on the definitions of one-to-oneand onto, and derive a formula for f−1. If no then explain why not. I did: assume f(m1,n1) = f(m2,n2) and I just equated both side of the equation making a sys. of equations and finding that m1 = m2 and n1=n2 making it one to one. However I can't prove how it is onto, any help?

OpenStudy (anonymous):

It seems more complicating cuz its a function that is dependent on 2 variables m and n Lemme just think how we would go abt this

OpenStudy (anonymous):

I know I can replace a and b in the equation, but seems kinda hard to evaluate

OpenStudy (anonymous):

Ok i think i got smth

OpenStudy (anonymous):

I am getting thrown off cuz its a multivariate function ... Let set up our function as follows: 3m+7n=a 2m+5n=b $$\frac{a-3m}{7}=n$$ So $$2m+5n=b \to 2m+5\left( \frac{a-3m}{7} \right )=b$$ $$2m+\frac{5}{7}a-\frac{3}{7}m=b$$ $${11}{7}m=b-\frac{5}{7}a$$ $$m=\frac{7b-5a}{11}$$

OpenStudy (anonymous):

but, how does proving it in a system of equation prove that it is onto

OpenStudy (anonymous):

Hey gimme a sec ... lemme think abt this

OpenStudy (misty1212):

HI!@!

OpenStudy (anonymous):

Ok well the way im thinking abt it is as follows: We basically have 2 equations for each coordinate f(m,n)=(a,b) Now f(m,n)=3m+7n=a for the x-coordinate and f(m,n)=2m+5n=b for the y-coordinate Now obv we can rewrite one of the equations in terms of n Once we have one of the equations in terms of n we can replace the n in the other equation since thats what n equals And then we can solve the newly revised equation in terms of m

OpenStudy (misty1212):

pick an arbitrary element of \(\mathbb{Z}\times \mathbb{Z}\) say \((p,q)\) and show that there is something that gets mapped to it it is the same as finding the inverse exactly

OpenStudy (anonymous):

Ya but its difficult finding the inverse when its a multivariate function

OpenStudy (anonymous):

Therefore the way to go about this is setting the coordinates as a and b And then use one equation to eliminate one variable

OpenStudy (anonymous):

Once we have done that then its easy to find its inverse

OpenStudy (anonymous):

Wait, so once you eliminated m ( or n) what would you do afterwards? plug it back in in the function to find its inverse?

OpenStudy (anonymous):

Reread My original post ... I first solved for n in my first equation Then I plugged that in to my second equation and eliminated the n And then I found my inverse And my inverse's domain was \(Y \in R\)

OpenStudy (anonymous):

hence it is onto

OpenStudy (anonymous):

completely sorry, my browser is screwing up some part of openstudy. I just saw, but your inverse is in terms of a,b so it is for some arbitrary value. Does that change anything? I know proving there is an inverse ( so finding a function in term of m,n) would prove its bijection.

OpenStudy (anonymous):

ya thats no problem We are choosing an arbitrary term for f(m,n) $$(a,b) \in f(m,n)$$

OpenStudy (anonymous):

Basically (a,b) is an element of the range

OpenStudy (anonymous):

ah so a,b is a range for which there exist a function to describe m. but if you can prove it for a smaller set, It also holds true for the whole domain?

OpenStudy (anonymous):

hmmmm I wldnt say a,b is the range but rather (a,b) is an arbitrary point in the range or in other words f(m,n) So basically are dependent values are m and n We give m and n and then according to what we set m and n as that's what our range , a and b , will be So basically view f(m,n) as our "y" Now when we are proving that a function is onto we are proving that every y in our codomain has an x-value associated So we find our inverse and show that if we plug in ANY SINGLE VALUE for a or b - which is our range - we will definitely find our m and n For example \(y=x^2-1\) is not onto So lets find our inverse which is : $$x=\sqrt{y+1}$$ Now for y<1 this function does not exist and therefore the function is not onto So we can basically bring the same principle to this multivariate example

OpenStudy (anonymous):

thanks a lot , you were super helpful!!!